how many pounds of candy worth $4.50/lb must be mixed with 100 pounds worth $3.50/lb to produce a mixture to sell for $4.00/lb
Answer by checkley77(12844)

4.50x+3.50*100=4(100+x)
4.50x+350=400+4x
4.50x-4x=400-350
.5x=50
x=50/.5
x=100 pounds of $4.50 candy is used.
Proof:
4.50*100+3.50*100=4(100+100)
450+350=4*200
800=4*200
800=800